40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 6 22:47:41 2021, gautam, Update, ALS, Noisy ALS ALSX_excess.pngbudget.pdf
    Reply  Thu Jan 7 19:16:11 2021, gautam, Update, ALS, Noisy ALS ALS_ERR_MON.pdfAUXnoise.pdf
       Reply  Thu Jan 7 20:07:27 2021, Koji, Update, ALS, Noisy ALS 
          Reply  Thu Jan 7 21:16:22 2021, gautam, Update, ALS, Noisy ALS 
             Reply  Thu Jan 7 23:25:19 2021, Koji, Update, ALS, Noisy ALS 
          Reply  Fri Jan 8 20:01:11 2021, gautam, Update, ALS, Noisy ALS ALSX_IR_green.pdf
Message ID: 15756     Entry time: Fri Jan 8 20:01:11 2021     In reply to: 15753
Author: gautam 
Type: Update 
Category: ALS 
Subject: Noisy ALS 

I did this test today. The excess noise around 100 Hz doesn't show up in the green beat.

See Attachment #1. The setup was as usual:

  • X-Arm cavity length stabilized to PSL frequency using the POX locking loop.
  • EX laser frequency locked to the X-Arm cavity length using the AUX PDH loop.
  • The "BEATX" channel records frequency fluctuations in the beat sensed on the IR beat photodiode, while the "BEATY" channel records frequency fluctuations in the beat sensed on the Green beat photodiode.
  • Since the green beat frequency fluctuations are twice that of the IR beat frequency fluctuations, I scaled the former ASD by a factor of 0.5 so as to compare apples to apples.
  • At low frequencies, the green beat is noisier, but that channel doesn't show the excess noise at mid frequencies you see in the IR beat. So the AUX PDH sensing noise is not to blame I think.

So, this excess appears to truly be excess phase noise on the fiber (though I have no idea what the actual mechanicsm could be or what changed between Aug and Oct 2020 that could explain it. Maybe the HEPA?

For this work, I had to spend some time aligning the two green beams onto the beat photodiode. During this time, I shuttered the PSL, disabled feedback via the FSS servo, turned the HEPA high, and kept the EX green locked to the arm so as to have a somewhat stable beat signal I could maximize. Everything has been returned to nominal settings now (obviously, since I locked the arms to get the data).


You may ask, why do we care. In terms of RMS frequency noise, it would appear that this excess shouldn't matter. But in all my trials so far, I've been unable to transition control of the arm cavity lengths from POX/POY to ALS. I suppose we could try using the green beat, but that has excess low frequency noise (which was the whole point of the fiber coupled setup). 

Quote:

How about resurrecting the PSL table green beat for the X arm to see if the non-fiber setup shows the same level of the freq noise (e.g. the PDH locking became super noisy due to misalignment etc).

Attachment 1: ALSX_IR_green.pdf  47 kB  | Hide | Hide all
ALSX_IR_green.pdf
ELOG V3.1.3-